16) Ethanol, CHsO, is produced by the fermentation of sugars in wine and beer. Given that ethanol has a molecular weight of 46.07 g/mol and a density of 0.789 g/em answer the following questions: a) How many mL of ethanol are consumed in a 12 oz. glass of red wine that contains 13.4 % ABV? Note: ABV stands for alcohol (ethanol) by volume. You must use the following conversion factors in your calculation: 128 oz- 1 gal; 1 gal 3.785 L b) How many molecules of ethanol would be in this glass of wine? c) What is the molarity of ethanol in the 12 oz glass of wine?
d) After pouring your 12 oz glass of wine, you cover the top so it -stays good for later. Wine also other molecules known as polyphenols that impart flavor, and are typically found at concentrations 2.568 grams per liter of wine. Wine tanins, for instance, are polyphenols that give wine a dry taste. Assuming all of the polyphenols in your wine are the tannin shown below, which are non-volatile, what would is the vapor pressure just due to ethanol in your sealed glass of wine? Pure ethanol has a vapor pressure of 7.869 KPa at room temperature. (Dont forget, you also have water in your glass of wine! Assume a density of 1.00 g/mL for water at room temp) C4H2s02 e) What would be the vapor pressure just due to water in your glass of wine? Pure water has a vapor pressure of 3.17 KPa at room temperature. What is the total vapor pressure of the solution? 2) You decide to take your sealed glass of wine to school and measure the vapor pressure of your solution to determine if your chemistry professor actually knows what hes talking about in class. You hook a labquest pressure sensor onto the top of the lid and measure the vapor pressure of the wine to this compare to your calculated value? Is this an ideal solution? What does this tell you about the intermolecular forces in your wine? Explain your Reasoning. be 24 torr. How does
Answer #1

a)

Volume of 12oz glass = 12 oz * 1 gal /128 oz * 3.785 L / 1 gal

= 0.35584 L = 355.84 mL

Volume of ethanol = 13.4 * 355.84 / 100

= 47.55 mL

b)

Mass of ethanol = Volume of ethanol * Density of ethanol

= 47.55 mL * 0.789 g /mL

= 37.5 g

Moles of ethanol present = Mass of ethanol / MW of ethanol

= 37.5 g / 46.07 g/mol

= 0.81

Molecules of ethanol = NA * Moles of ethanol present

= 6.022 x 1023 * 0.81

= 4.9 x 1023

c)

Molarity of ethanol = Moles of ethanol / Total Volume (L)

= 0.81 / 0.35584 L

= 2.3 M

d)

Mass of polyphenols = 0.35584 L * 2.568 g / L

= 0.91 g

MW of C34H28O22 = 34 * 12 + 1 * 28 + 16 * 22 = 788 g/mol

Moles of polyphenols = 0.91 / MW of C34H28O22

= 0.00116

Volume of water in wine glass = 355.84 mL - 47.55 mL

= 307.3 mL

Mass of water in wine glass = 307.3 mL * 1 g/mL = 307.3 g

Moles of water in wine glass = 307.3 g / 18 g/mol

= 17.07

Mole fraction of ethanol, X = Moles of ethanol / Total moles of all components

= 0.81 / (0.81 + 0.00116 + 17.07)

= 0.0455

Pure ethanol vapor pressure, P0 = 7.869 kPa

Vapor pressure just due to ethanol = P0 * X

= 7.869 kPa * 0.0455

= 0.36 kPa